assum that Z-scores are normally distrubuted with a mean of 0 and a standard deviation of 1.
Find p(Z more than or equal to 1.45)

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi
Find p(Z more than or equal to 1.45) = .0735

Answer by stanbon(75887)

If IQ scores are normally distributed with a mean of 100 and a standard deviation of 15.
- Find the probability that the mean IQ score is between 90 and 95, if 9 IQ scores are randomly selected.
---
t(90) = (90-100)/[15/sqrt(100)] = -10/[1.5] = -6.6667
t(95) = (95-100)/[15/sqrt(100)] = -5/[1.5] = -3.3333
-----
P(90< x-bar <95) = P(-6.6667< t < -3.3333 when df=8) = 0.0051
====================
Cheers,
Stan H.
====================